A box contains about 6.68 x 10^21 hydrogen atoms at room temperature 21 degrees C find the thermal energy of these atoms answers in J
Answer #1

Solution :

Given :

N = 6.68 x 1021

T = 21oC = 294.15 K

Since thermal energy is given by : E = (3/2) N kB T

Where, kB = 1.3807 x 10-23 J/K.

Thus : E = (3/2)(6.68 x 1021)(1.3807 x 10-23 J/K)(294.15 K) = 40.69 J